# Zero-Shot LLMs in Human-in-the-Loop RL: Replacing Human Feedback for Reward Shaping11footnotemark: 1

### v1

@@ A.2. Custom Edge Case Scenarios:
    Scenario 1: Congested Lane
    The rightmost lane, typically rewarded in highway-v0, is heavily congested with slower-moving vehicles (speeds between 15-20 m/s). The adjacent lanes are moderately occupied with faster-moving traffic (25-30 m/s). The ego vehicle starts in the rightmost lane at 25 m/s and must navigate around congestion to maintain high speed while avoiding collisions. The scenario tests the feedback system’s ability to balance rewards for staying in the right lane against necessary lane changes. The duration parameter in the highway environment determines the maximum number of timestep an episode can run (if the termination condition is not met) is set to 40 by default.
@@
    Scenario 2: Slow-Moving Obstacle in Middle Lane
@@ 5.2 Metrics Used
    where $R_{C_{e}}$ is the cumulative reward for episode $e$, $E$ is the total number of episodes,$T_{e}$ is the number of timesteps in episode $e$ and $R_{e,t}$ is the reward at timestep $t$ in episode $e$.
@@ 5.3 Results and Discussion
FIS FMA ATT AER
HF-RSM 0.0 40.0 28.472
@@
biased HF-RSM (AGG) 0.4 32.8 24.269
biased HF-D (AGG) 6.0 10.2 9.006
biased HF-RSM(CON) 0.4 32.8 24.269
biased HF-D(CON) 5.6 8.2 7.039
@@
Table 4: Performance of different FIS, on default Highway Environment
Our experimental results ( see Table4 ) clearly demonstrate that even minor biases in human feedback can substantially degrade reinforcement learning agent performance. In the default highway environment, the unbiased approaches (HF-RSM, HF-D, LLM-D, LLM-HFBF) consistently achieved optimal performance with an Average Episodic Reward (AER) of 28.472, zero Feedback Misalignment (FMA), and maximum Average Terminate Time (ATT) of 40.0. In contrast, biased human feedback led to significant performance deterioration. The biased HF-D approaches showed the most severe performance drops, with aggressive bias (AGG) resulting in AER dropping to 9.006, FMA increasing to 6.0, and ATT falling to 10.2, while conservative bias (CON) caused AER to drop further to 7.039, FMA to reach 5.6, and ATT to decline to 8.2. This suggests that direct biased human feedback introduces substantial misalignment with optimal driving behavior, resulting in more frequent collisions and shorter episode durations.
The Human Feedback with Reward Shaping Model (HF-RSM) demonstrated greater resilience to bias compared to direct human feedback (HF-D). Even when biased, HF-RSM maintained relatively better performance with an AER of 24.269, lower FMA of 0.4, and higher ATT of 32.8. This indicates that surrogate models can partially mitigate the effects of bias through their generalization capabilities, acting as a buffer between potentially skewed human evaluations and the agent’s learning process. The LLM-based direct feedback approach (LLM-D) exhibited remarkable effectiveness, achieving optimal performance metrics (AER: 28.472, FMA: 0, ATT: 40.0) equivalent to unbiased human feedback. This validates the hypothesis that large language models can effectively replace human feedback without requiring task-specific training. Despite having no prior specialized training on highway navigation tasks, the LLM provided feedback comparable to unbiased human input.
FIS FMA ATT AER
Case 1 HF-D 5.0 20.2 14.174
biased HF-D (AGG) 6.4 8.2 7.122
biased HF-D (CON) 6.8 8.4 7.187
@@
Case 2 HF-D 5.0 19.6 13.299
biased HF-D (AGG) 6.6 7.8 6.839
biased HF-D (CON) 6.4 4.6 3.596
@@
The edge case experiments in Table 5 further highlight the robustness of LLM-based approaches. In both challenging scenarios, LLM-D matched the performance of unbiased HF-D (AER: 14.174 in Case 1, 13.299 in Case 2), while biased feedback approaches continued to produce suboptimal results. In Case 1, biased HF-D with aggressive bias (AGG) had an AER of 7.122 with FMA of 6.4, while conservative bias (CON) had an AER of 7.187 with FMA of 6.8. In the more challenging Case 2, biased HF-D performed even worse, with aggressive bias resulting in an AER of 6.839 and FMA of 6.6, while conservative bias led to the lowest performance with an AER of just 3.596 and FMA of 6.4. This demonstrates how rule-based approaches struggle with contextual adaptation, particularly when optimal behavior requires deviating from rigid predefined rules.
These findings have several important implications for Human-In-The-Loop Reinforcement Learning systems. First, systems incorporating human feedback should explicitly account for potential biases through detection and mitigation mechanisms to prevent performance degradation. Second, modern large language models possess sufficient contextual understanding to provide effective feedback for reinforcement learning tasks without requiring domain-specific training, enabling more scalable RL systems. Third, combining human expertise with LLM-based bias detection appears to offer the most promising path forward, allowing for selective human intervention only when necessary while maintaining optimal performance. Finally, the successful application of LLM-based feedback in continuous control tasks suggests broader applicability beyond NLP domains, potentially extending to various other reinforcement learning domains.
@@ 6 Limitations and Future work
Despite promising results, several important limitations warrant further investigation. First, our current implementation requires trajectory data frames to be relatively small due to LLM context window constraints, limiting the historical information available for feedback generation. Second, obtaining feedback from LLMs incurs significant computational costs, requiring GPU resources that may not be available in all application contexts, especially for real-time decision-making.
Third, our approach depends heavily on iterative and precise prompt design, which requires expertise and may need adjustment across different environments. Additional limitations include environmental complexity beyond highway navigation, the need to explore different LLM architectures, and methods for real-time LLM-based feedback. Future work will address these limitations and explore applications to a broader range of continuous control tasks, including robotic manipulation and autonomous driving in diverse scenarios.

### v2

@@ A.2. Custom Edge Case Scenarios:
    Case 1: Congested Lane
    The rightmost lane, typically rewarded in highway-v0, is heavily congested with slower-moving vehicles (speeds between 15-20 m/s). The adjacent lanes are moderately occupied with faster-moving traffic (25-30 m/s). The ego vehicle starts in the rightmost lane at 25 m/s and must navigate around congestion to maintain high speed while avoiding collisions. The highway environment’s duration parameter, which sets the max timesteps per episode if not terminated, is set to 40 by default.
@@
    Case 2: Slow-Moving Obstacle in Middle Lane
@@
A.2. Reacher (Reacher-v5)
Reacher-v5 in Gymnasium is a robotic control environment where a two-jointed arm must reach a target in a 2D plane. The state space is 11-dimensional, including joint angles, angular velocities, and relative positions of the target and end-effector. The action space is continuous (2D), representing torques applied to the joints. The reward function is given as
$r=-w_{\text{near}}\|p_{\text{eff}}-p_{\text{tgt}}\|_{2}-w_{\text{ctrl}}\|a\|_{2}^{2}$
where $p_{eff}$ (end-effector position) is the 2D coordinate of the fingertip of the robotic arm, $p_{tgt}$ (target position) is the 2D coordinate of the goal location, and $\|p_{eff}-p_{tgt}\|_{2}$ is the Euclidean distance between them, which the reward function minimizes to encourage precise reaching, $w_{\text{near}}=1.0$, $w_{\text{ctrl}}=0.01$. Encourages target proximity, penalizes large actions.
@@ 5.2 Metrics Used
[…]
@@ 5.3 Results and Discussion
Our experimental results (see Table 4) clearly demonstrate that even minor biases in human feedback can substantially degrade the performance of the reinforcement learning agent. In the default highway environment, the unbiased approaches (HF-RSM, HF-D, LLM-D, LLM-HFBF) consistently achieved near-optimal performance. The Average Episodic Reward (AER) for these methods ranged from 27.608 to 28.472. The Feedback Misalignment (FMA) was 0.6 for HF-RSM and 0.0 for HF-D, LLM-D, and LLM-HFBF. All of these unbiased methods achieved a maximum Average Terminate Time (ATT) of 40.0.
In contrast, biased human feedback led to significant performance deterioration. The biased HF-D approaches showed the most severe performance drops. With aggressive bias (AGG), the AER dropped to 5.863, FMA increased to 5.6, and ATT fell to 7.0. Reckless adaptive bias (RAD) caused the AER to drop further to 1.840, FMA to reach 5.0, and ATT to decline to 3.0. This suggests that direct biased human feedback introduces substantial misalignment with optimal driving behavior, resulting in more frequent collisions and shorter episode durations. .
FIS FMA ATT AER (Highway)
HF-RSM 0.6 40.0 27.608
@@
biased HF-RSM (AGG) 0.6 40.0 29.431
biased HF-D (AGG) 5.6 7.0 5.863
biased HF-RSM(RAD) 1.6 34.4 25.448
biased HF-D(RAD) 5.0 3.0 1.840
@@
Table 4: Performance of different FIS, on default Highway Environment. The policy is trained using PPO[26] from StableBaselines3 (SB3) implementation[32], for 0.01 Million timesteps.
The Human Feedback with Reward Shaping Model (HF-RSM) demonstrated greater resilience to bias compared to direct human feedback (HF-D). Even when biased, HF-RSM maintained relatively better performance with an AER of 25.448, lower FMA of 1.6, and higher ATT of 34.4. This indicates that surrogate models can partially mitigate the effects of bias through their generalization capabilities, acting as a buffer between potentially skewed human evaluations and the agent’s learning process. The LLM-based direct feedback approach (LLM-D) exhibited remarkable effectiveness, achieving optimal performance metrics (AER: 28.472, FMA: 0, ATT: 40.0) equivalent to unbiased human feedback. This validates the hypothesis that large language models can effectively replace human feedback without requiring task-specific training. Despite having no prior specialized training on highway navigation tasks, the LLM provided feedback comparable to unbiased human input.
FIS FMA ATT AER (Highway)
Case 1
HF-D 5.0 20.2 14.174
biased HF-D (AGG) 5.8 5.4 4.332
biased HF-D (RAD) 8.8 7.4 5.856
@@
Case 2
HF-D 5.0 19.6 13.299
biased HF-D (AGG) 6.6 7.8 6.812
biased HF-D (RAD) 8.0 6.0 4.682
@@
The edge case experiments in Table 5 further highlight the robustness of LLM-based approaches. In both challenging scenarios, LLM-D matched the performance of unbiased HF-D (AER: 14.174 in Case 1, 13.299 in Case 2), while biased feedback approaches continued to produce suboptimal results. In Case 1, biased HF-D with aggressive bias (AGG) had an AER of 4.332 with FMA of 5.8, while reckless adaptive bias (RAD) had an AER of 5.856 with FMA of 8.8. In the more challenging Case 2, biased HF-D again performed poorly, with aggressive bias resulting in an AER of 6.812 and FMA of 6.6, while reckless adaptive bias led to the lowest performance with an AER of just 4.682 and FMA of 8.0. This demonstrates how rule-based approaches struggle with contextual adaptation, particularly when optimal behavior requires deviating from rigid predefined rules.
FIS AER (Reacher)
HF-D -38.210
biased HF-D (AGG) -80.336
biased HF-D (RAD) -30.009
LLM-D -11.637
LLM-HFBF -12.164
Table 6: Comparison of performances of different FIS on the Reacher environment (The policy is trained using PPO [26] from SB3 [32] implementation, for 0.2 Million time-steps). The HF-D biased and unbiased rules and the LLM prompt structure are conceptually similar to that of the Highway case as elaborated in Sec. 4.1.3.
To further assess the robustness and generalizability of LLM-based approaches, we tested them in a second continuous control environment, Reacher. The results, shown in Table 6, reinforce the trend observed in the Highway environment. LLM-based approaches continued to match the performance of unbiased HF-D, while biased feedback approaches resulted in substantial performance degradation.
[…]
Figure 3: Performance Degradation Due to Bias Across Environments: compares the performance degradation of different feedback strategies (LLM-D, LLM-HFBF, and two biased HF-D variants) relative to the unbiased human feedback (HF-D) baseline. The values on the y-axis represent the percentage decrease from the optimal performance achieved by HF-D in each environment, with lower percentages indicating less degradation. The data from the Highway and Reacher environments are shown side-by-side. The optimal baseline Average Episodic Reward (AER) for these calculations is 28.472 for Highway and -11.637 for Reacher. The final values are normalized to a 0–100 scale using Min–Max normalization
These findings have key implications for Human-In-The-Loop Reinforcement Learning. First, detecting and mitigating human biases is crucial to maintaining optimal performance. Second, large language models can provide effective feedback without domain-specific training, enabling more scalable RL systems. Third, integrating human expertise with LLM-driven bias detection offers a balanced approach, allowing selective human intervention only when needed. Finally, the success of LLM-based feedback in continuous control tasks suggests its broader applicability across diverse RL domains.
6 Conclusion
In this work we successfully demonstrate that zero-shot LLMs can effectively replace human feedback for reward shaping in continuous control tasks. Through experiments, we showed that biased human feedback can significantly degrade RL agent performance, with reductions of up to 94% in average episodic reward. In contrast, the proposed LLM-based methods (LLM-D and LLM-HFBF) maintain performance levels comparable to unbiased feedback, even in challenging edge-case scenarios. Moreover the LLM-HFBF hybrid framework offers a promising solution that addresses both the inherent biases in human supervision and the limitations of LLMs, creating a more balanced and reliable system for HITL-RL applications.
@@ 7 Limitations and Future work
Despite promising results, several important limitations warrant further investigation. First, our current implementation requires trajectory data frames to be relatively small due to LLM context window constraints, limiting the historical information available for feedback generation. Second, obtaining feedback from LLMs incurs significant computational costs, requiring GPU resources that may not be available in all application contexts, especially for real-time decision-making. Third, our approach depends heavily on iterative and precise prompt design, which requires expertise and may need adjustment across different environments. Additional limitations include environmental complexity beyond highway navigation, the need to explore different LLM architectures, and methods for real-time LLM-based feedback. Future work will address these limitations and explore applications to a broader range of continuous control tasks, including robotic manipulation and autonomous driving in diverse scenarios.